What volume of 0.100 M K2S solution contains 3.5 g of the solute?


1
Expert's answer
2022-05-18T14:10:02-0400

1. From the formula CM = n / V we derive the formula for finding V (volume):

V = n / CM .

V– volume

n– amount of substance ( mol )

CM– concentration molarity

2. We find the amount of substance (n). To do this, we divide the mass of K2S (110g/mol) by its molecular mass:

n=m/M = 3.5 / 110 = 0.03182 mol

3. We find the volume using the formula V=n/CM:

V = n / CM = 0.03182 / 0.1 = 0.3182 litr.

ANSWER: 0.3182 litr or 318.2 ml